OpenStudy (anonymous):

What is the y-value of the y-intercept of the line 3x + 4y = 16?

OpenStudy (anonymous):

Writing the eq 3x + 4y = 16 in y = mx +c form we find that 4y = -3x + 16 y = (-3/4)x + 16/4 y = (-3/4)x + 4 comparing it with y = mx +c we find that c=4 which is the value of the y- intercept